OpenStudy (anonymous):

Essay: Show all work. A gardener wants to create a rectangular garden with length of 2x-3y ft. and width of x+4y ft. What is an algebraic expression for the area of the garden? Be sure to multiply this out, express in simplest correct mathematical form, and include units. Walk me through, please? (2x-3y)(x+4y) is this correct?

OpenStudy (anonymous):

F=2x*x = 2x^2 O=2x*4y= 8xy I=-3y*x= -3xy L=-3y*4y =-12y^2 =2x^2+8xy-3xy-12y^2 =2x^2+5xy-12y^2

OpenStudy (anonymous):

use the FOIL method to multiply the two First- (2x)(x)=2x^2 Outside- (2x)(4y)=8xy Inside- (-3y)(x)=-3xy Last- (-3y)(4y)=-12y Combining all the terms: 2x^2+8xy-3xy-12y^2 Simplifying: 2x^2+5xy-12y^2

OpenStudy (anonymous):

Thank you for your help. I wasn't for sure if the xy thing was right or not.

OpenStudy (anonymous):

Do I need to also put in there ft^2?